a student weighed a clean and dry beaker and found its mass to be 53.6 g. She then poured chloroform in the beaker and then measured the total mass to be 74.8 g. Help the student identify the volume of the liquid poured into the beaker


1
Expert's answer
2021-02-09T03:39:32-0500

"m_{chl} = 74.8 -53.6 = 21.2 \; g"

We know that the density ρ of chloroform is 1.49 g/mL.

"ρ = \frac{m}{V} \\

V = \frac{m}{ρ} \\

= \frac{21.2}{1.49} \\

= 14.23 \;mL"

Answer: 14.23 mL